### v1

@@ 4.1 Results
Figure 2 shows the curves for comparing the algorithms on the MNIST dataset. The S+t-SNE algorithm is represented by the continuous line and the t-SNE by the dashed line. Each point in the line represents an action by the algorithm, typically at the end of the batch. The main point of comparison is the initial slice of the data used to define the opening space. That is the quantity of data accumulated before projecting the first time. The slice is represented fractionally regarding the total amount of data from 00 to $1$. The lines representing S+t-SNE represent the variation in the size of batch (B), the number of PEDRUL (or density points) considered (D), e.g. D=300, the projection holds only 300 PEDRUL at all times, and the number of iterations allowed for the algorithm to run (Iter).
@@
As expected, the peak memory used by t-SNE (2) aggressively increases as the number of points increases, achieving the same memory peak in all slices in the last reprojection. As for S+t-SNE, the peak memory use is achieved when getting the opening projection. The initial S+t-SNE iteration (getting the opening projections) for slice $x$ uses the same process of a t-SNE application for that slice $x$. However, the peak memory of S+t-SNE is higher because it has to account for the search of the PEDRUL points. We notice that larger slices further increase the initial memory peak gap between algorithms because searching for PEDRUL points within more data is more costly. After the first iteration, the peak memory reduces drastically and remains constant.
@@
The results for execution time (2) were as expected. Larger batches, number of points and iterations increase the computational time. Furthermore, a large slice increases the initial time of S+t-SNE but contributes to a faster decrease in the time consumed.
Figure 3: Projections by S+t-SNE for MNSIT. B:400 D:400 IT:700 Slice:0.2
@@
Figure 3 shows the evolution of the S+t-SNE embeddings for MNIST for B:400 D:400 IT:700 Slice:0.2. Besides the points delimiting the hulls, the PEDRUL are also visible. All ten classes are well distinguishable. Moreover, ECS eliminates inaccurately projected points, as posited in section 3.1. One example is in the transition from the top-right figure to the bottom-left, where the blue cluster had its stretched side removed. A typical ECS cut.
@@
From figure 4, we can see that increasing the batch size and number of points increases the peak memory consumption. However, the increase is very small (in the order of single bytes), even though we are doubling batch size and number of points. This result happens possibly due to the capacity of larger batches and the larger number of points, allowing us to pick points with better representational capabilities, lasting more iterations and triggering fewer operations and consequently less accumulation of objects in memory. The results confirm the ones obtained with MNIST: the memory used is near constant after reaching a stable point where the representation is stable (initial peak removed in image).
@@
We believe the considerations taken from MNIST regarding the value to use for each parameter hold in this case. Figure 5 shows the projections for the synthetic data. The bottom-right plot shows the embeddings if no ECS was used. The plots show that ECS allows us to focus on the points of interest. Hence, we can see the most recent movements of the dataset, being the gradual shrinking of the green structure and the orange and blue ones slowly drifting closer to each other until they overlap.
The effect of ECS is not noticeable in the computational time due to its small computational time footprint. Even without a parallel implementation, the main difference between configurations is the number of hulls/clusters found at each iteration. However, this number is similar to all configurations because the clustering parameters and dataset are the same for all configurations. Furthermore, the number of PEDRUL points has a constant effect in the time taken by ECS, and since the gap between configurations in this regard is not very large the effect is negligible. The drift has virtually no impact on memory consumption because of its gradual nature.
@@ 5 Conclusion
In this paper, we developed an efficient adaptation of t-SNE called S+t-SNE to work with data streams. Our version supports dimensionality reduction of online data and can detect drift. A possible direction for future work is to test different mechanisms for obtaining PEDRUL points and test the robustness of the method to different types of drift. In order to allow for a uniform comparison across online methods, it would be interesting to develop a metric for the comparison of projections of arbitrary size.

### v3

@@ 4.1 Results
Figure 2 shows the curves for comparing the algorithms on the MNIST dataset. The continuous lines and the t-SNE by the dashed lines represent the S+t-SNE algorithm. Each point in a line represents an action by the algorithm, typically at the end of the batch. The main point of comparison is the initial slice of the data used to define the opening space. That is the quantity of data accumulated before projecting the first time. The slice is represented fractionally regarding the total amount of data from 00 to $1$. The lines representing S+t-SNE represent the variation in the size of batch (B), the number of PEDRUL (or density points) considered (D), and the number of iterations allowed the algorithm to run (Iter).
@@
As expected, the peak memory used by t-SNE (figure 2) aggressively increases as the number of points increases, achieving the same memory peak in all slices in the last reprojection. As for S+t-SNE, the peak memory use is achieved when getting the opening projection. The initial S+t-SNE iteration (getting the opening projections) for slice $x$ uses the same process of a t-SNE application for the same slice $x$. However, the peak memory of S+t-SNE is higher because it has to account for the search of the PEDRUL points. We notice that larger slices further increase the initial memory peak gap between algorithms because searching for PEDRUL points within more data is more costly. After the first iteration, the peak memory reduces drastically and remains constant.
@@
The results for execution time (figure 2) were as expected. Larger batches, number of points and iterations increase the computational time. Furthermore, a large slice increases the initial time of S+t-SNE but contributes to a faster decrease in the time consumed.
Figure 3: Projections by S+t-SNE for MNIST. B:400 D:400 IT:700 Slice:0.2
@@
Figure 3 shows the evolution of the S+t-SNE embeddings for MNIST for B:400 D:400 IT:700 Slice:0.2. Besides the points delimiting the hulls, the PEDRUL are also visible. All ten classes are well distinguishable. Moreover, ECS eliminates inaccurately projected points, as posited in section 3.1. One example is the transition from the top-right figure to the bottom-left, where the blue cluster had its stretched side removed. It’s a typical ECS cut.
@@
From figure 4, we can see that increasing the batch size and number of points increases the peak memory consumption. However, the increase is very small (in the order of single bytes), even though we are doubling batch size and number of points. This result happens due to the capacity of larger batches and the larger number of points, allowing us to pick points with better representational capabilities, lasting more iterations and triggering fewer operations and consequently less accumulation of objects in memory. The results confirm the ones obtained with MNIST: the memory used is near constant after reaching a stable point where the representation is stable (initial peak removed in image).
@@
We believe the considerations taken from MNIST regarding the value to use for each parameter hold in this case. Figure 5 shows the projections for the synthetic data. The bottom-right plot shows the embeddings if no ECS was used and all points were considered as PEDRUL. The plots show that our techniques allow us to focus on the points of interest. Hence, we can see the most recent movements of the dataset, these being the gradual shrinking of the green structure and the orange and blue ones slowly drifting closer to each other until they overlap.
The effect of ECS is not noticeable in the computational time due to its small runtime footprint. Even without a parallel implementation, the main difference between configurations is the number of hulls/clusters found at each iteration. However, this number is similar to all configurations because the clustering parameters and dataset are the same for all configurations. Furthermore, the number of PEDRUL has a constant effect on the time taken by ECS, and since the gap between configurations in this regard is not very large, the effect is negligible. The drift has virtually no impact on memory consumption because of its gradual nature.
@@ 5 Conclusion
In this paper, we developed an efficient adaptation of t-SNE called S+t-SNE to work with data streams. Our version supports dimensionality reduction of online data and can adapt to data drift. A possible direction for future work is to test different mechanisms for obtaining PEDRUL and test the robustness of the method to different types of drift. To allow for a uniform comparison across online methods, it would be interesting to develop a metric for the comparison of projections of arbitrary size.